# Portrait3D: 3D Head Generation from Single In-the-wild Portrait Image

### v1

@@ Abstract
While recent works have achieved great success on one-shot 3D common object generation, high quality and fidelity 3D head generation from a single image remains a great challenge. Previous text-based methods for generating 3D heads were limited by text descriptions and image-based methods struggled to produce high-quality head geometry. To handle this challenging problem, we propose a novel framework, Portrait3D, to generate high-quality 3D heads while preserving their identities. Our work incorporates the identity information of the portrait image into three parts: 1) geometry initialization, 2) geometry sculpting, and 3) texture generation stages. Given a reference portrait image, we first align the identity features with text features to realize ID-aware guidance enhancement, which contains the control signals representing the face information. We then use the canny map, ID features of the portrait image, and a pre-trained text-to-normal/depth diffusion model to generate ID-aware geometry supervision, and 3D-GAN inversion is employed to generate ID-aware geometry initialization. Furthermore, with the ability to inject identity information into 3D head generation, we use ID-aware guidance to calculate ID-aware Score Distillation (ISD) for geometry sculpting. For texture generation, we adopt the ID Consistent Texture Inpainting and Refinement which progressively expands the view for texture inpainting to obtain an initialization UV texture map. We then use the id-aware guidance to provide image-level supervision for noisy multi-view images to obtain a refined texture map. Extensive experiments demonstrate that we can generate high-quality 3D heads with accurate geometry and texture from a single in-the-wild portrait image. The project page is at https://jinkun-hao.github.io/Portrait3D/.
Figure 1: Given an in-the-wild portrait image, Portrait3D has the ability to generate high-quality 3D head with detailed face geometry and photo-realistic texture. With stylized diffusion models, Portrait3D can be used to customize 3D heads with multiple styles while keeping the identity in the portrait image.
@@ 1 Introduction
3D head generation aims at creating high-quality 3D digital assets of human heads that align with user preferences utilizing various input data, which has significant application in film character creation, gaming, online meetings, education, etc. Conventional methods rely on multi-view geometric consistency [19, 51, 27, 46, 28] or statistical 3D face prior model [2, 52, 10] to achieve high-quality 3D head generation. The former necessitates multi-view images or several minutes of monocular video, which are not applicable to a single in-the-wild portrait image. Moreover, 3D face shape prior models typically focus on the face region, failing to accurately reconstruct the entire head shape and being unable to handle elements such as long hair and clothing. An alternative approach involves employing Generative Adversarial Networks (GANs) to accomplish 3D head generation, which is constrained by limited training data and typically produces frontal-facing faces. Moreover, they struggle to preserve identity consistency when presented with images of large head poses. Recently, text-to-image [32, 33, 37, 38, 14, 29] generation models trained on massive-scale image datasets have achieved significant success. Some approaches employ SDS loss [30] to ensure similarity between the generated 3D model and images, facilitating image-to-3D [26, 42, 24, 23, 25] or text-to-3D [30, 21, 39, 43, 44]. However, when employing these methods for customized 3D head generation, the absence of identity information makes it difficult to accurately reproduce the geometric details and texture realism of the head in the portrait image.
Through analyzing previous methods, we summarize the key challenges from the three stages in the optimization-based 3D head generation approach: 1) Geometry Initialization: Previous image-to-3D methods used an ellipsoid as the initial geometry [5] and could not accurately reconstruct the geometry of the head. On the other hand, text-to-head methods used a unified FLAME head [13, 22] as the geometric prior, which lacks personalized information and cannot generate individualized features that deviate from the FLAME head, such as long hair. 2) Geometry Sculpting: Although image-to-3D methods attempted to use view-dependent diffusion  [23]to guide geometry generation, the resulting geometry lacked the generalization ability to the detailed head geometry. Besides, text-to-3D methods use facial image description text to generate geometry, which can result in a deviation from the identity of the reference image. 3) Texture Generation: Previous methods used VSD loss [44] to mitigate color over-saturation issues based on SDS. However, the generated textures still tended to have excessive shading and unrealistic texture [22, 16].
To address the challenges mentioned above, we incorporate identity information into the three stages of 3D head generation compared with previous methods 2. Firstly, we extract Identity embedding using an ID feature extractor and align them with text embedding using a pre-trained ID adapter with decoupled cross-attention. This, along with landmark-guided ControlNet, enables ID diffusion and provides ID-aware guidance, allowing us to incorporate the identity and facial layout information from the reference image into the geometry and texture generation stages. Secondly, we propose ID-aware geometry supervision. By extracting canny maps containing detailed information from the portrait image, we combine canny-guided ControlNet and ID features with off-the-shelf text-to-normal and text-to-depth methods [16] to generate high-quality geometric constraint for the geometry generation stage. Additionally, we obtain ID-aware initialization using 3D GAN inversion. In the geometry sculpting stage, we optimize the 3D head using ID-aware guidance with ID-aware Score Distillation (ISD) loss. In the texture generation stage, we first use Progressive Texture Inpainting to generate a UV-textured map for the uncolored 3D head mesh. Then we leverage ID diffusion to denoise the rendered image from random viewpoints and further refine the texture by calculating the loss between the rendered and denoised images at the image level.
@@
In this paper, we propose Portrait3D, a new method for generating high-quality 3D heads from single in-the-wild portrait images. We present Identity-aware Head Guidance and Initialization (§4.1) and use ID-aware guidance to guide Geometry Generation (§4.2). In the texture generation stage, we propose ID Consistent Texture Inpainting and Refinement (§4.3), where we first use the inpainting approach to generate a rough texture and then use ID-aware guidance for texture refinement. With these methods, we can generate high-quality 3D head models with consistent identities from a single in-the-wild portrait image and further enable texture editing.
In summary, the contributions of our work are as follows:
@@
  • An ID Consistent Texture Inpainting and Refinement method, which uses Progressive Texture Inpainting and refinement approach to achieve photo-realistic texture generation.
2 Related work
@@
Recently, various approaches have explored reconstructing 3D head avatars from monocular face videos using different 3D representations. Among these approaches are mesh-based techniques like NHA [10], neural implicit representation such as NerFace [9], explicit point cloud methods like PointAvatar[52], and 3D Gaussian-based approaches like [45]. However, these methods necessitate several minutes of video input and are unable to reconstruct complete heads from single in-the-wild images. Moreover, they struggle to reconstruct less prominent parts in videos, such as the back of the head. In addition, generating heads from 3D GANs has rapidly developed. EG3D [3] proposed a hybrid tri-plane representation that can quickly and efficiently render high-quality facial images. Panohead [1] expanded on EG3D by increasing the number of planes, allowing for the generation of 3D faces from limited angles to the entire head. However, these methods are limited by the representation capability of implicit representation and cannot produce high-quality surface geometry.
With the development of text-to-image models, generating human heads directly from text has become an important application. Headsculpt [13] incorporates facial landmark information into the text-to-3D generation method, effectively constraining the facial layout and avoiding the Janus face problem. HeadArtist [22] builds on Headsculpt by proposing SSD loss and decoupling the geometry and texture generation processes, resulting in an even higher-quality human head generation. Additionally, Humannorm [16] fine-tunes the diffusion model on 3D human body datasets to obtain diffusion models for text-to-normal and text-to-depth, enhancing the generation ability of human body parts. However, the texture generated by these methods can be oversaturated, making it difficult to achieve photo-realistic head generation.
2.2 Text to 3D Generation
Recent advances in combining 3D generative architectures with image generative models make it possible to synthesize diverse high-fidelity 3D scenes and objects from text prompts. Efforts such as DreamFusion [30] explore text-to-3D generation by leveraging a Score Distillation Sampling (SDS) loss derived from a 2D text-to-image diffusion model and show impressive results. Subsequently, based on the promising SDS loss, Magic3D [21] increases the rendering resolution of generated 3D objects from 64 to 512. Fantasia3D [5] models the appearance via the BRDF modeling, generating compelling geometry and photorealistic object texture. ProlificDreamer [44] proposes Variational Score Distillation (VSD), an advanced guiding loss capable of solving oversaturation problems. Although many works have solved the problem of texture oversaturation to some extent, it is still difficult to generate realistic geometry and texture.
2.3 Image to 3D Generation
Image-driven 3D generation have also been greatly influenced by the emergence of diffusion models. Zero-1-to-3 first proposes view-dependent diffusion, which primarily focuses on learning camera viewpoints within large-scale diffusion models, enabling zero-shot novel view synthesis. GeNVS [4] leverages pixel-aligned features from input views to add 3D awareness to 2D diffusion models. On the other hand, some optimize-based image-to-3D methods, such as Magic123 [31] and Make-it-3D [42], incorporate 3D priors from Zero-1-to-3 and employ coarse-to-fine two-stage optimization. Make-it-3D, in particular, enhances texture and geometric structure in its fine stage, yielding high-quality textured point clouds as final outputs. Subsequent works [41, 48] have consistently aimed to build upon these achievements, further refining and enhancing the quality of generated 3D models. However, these methods are limited by the ability of view-dependent diffusion, which cannot accurately generate the geometry from the image.
3 Preliminaries
Score Distillation Sampling for 3D Generation. Recently, optimization based 3D generation approaches usually optimize a 3D scene using Score Distillation Sampling (SDS) [30] loss.
SDS loss is proposed to distillate 3D prior from 2D pretrained diffusion models by minimizing the rendering errors from an optimized 3D scene. Specifically, during optimization, an image $x_{0}=\mathcal{G}(\theta,c)$ is rendered from a 3D representation with parameter $\theta$ under the randomly sampled camera $c$ using a differentiable renderer $\mathcal{G}$. The SDS loss can be calculated as:
$\nabla_{\theta}L_{\mathrm{SDS}}=\mathbb{E}_{t,\boldsymbol{\epsilon}}[\omega(t)\|\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},t,y)-\boldsymbol{\epsilon}\|^{2}\frac{}{}],$ (1)
where $\omega(t)$ is a weighting function. $x_{t}$ is the noisy image by adding noise $\epsilon\sim\mathcal{N}(0,1)$ to rendered image $x_{0}$ at timestep $t$. $y$ is the conditioned text prompt. $\boldsymbol{\epsilon}_{\phi}(\cdot)$ is the denoising direction predicted by the pretrained diffusion model $\phi$. SDS encourages creating plausible 3D models that conform to the conditional distribution in diffusion latent space. To improve the geometry quality, the classic SDS loss uses a large weight of classifier-free guidance, which will result in high color saturation.
@@ 3 Preliminaries
Figure 3: Overview of our proposed Portrait3D framework. Given a single in-the-wild face image, Portrait3D uses Identity-aware Head Guidance and Initialization (§4.1) to integrate identity information into the Geometry Sculpting stage (§4.2). In the texture generation stage, ID Consistent Texture Inpainting and Refinement (§4.3) is applied to generate a high-quality head texture, where we first use the inpainting method to generate a rough texture and then use image-level ID-aware supervision for texture refinement. With these methods, we can generate high-quality 3D head models with consistent identities from a single in-the-wild face image.
Different Variations of SDS. Recent works [44, 31, 41] introduce different variations of SDS to enhance generation fidelity and training fidelity. Specifically, [44] proposed the Variational Score Distillation (VSD) to improve the texture and geometry quality. VSD treats the 3D optimization process as learning a 3D distribution rather than a single point under the conditional text prior. During the optimization, VSD relies on fine-tuned low-rank adaptation (LoRA) [15] model to represent the distribution of rendered images under different camera poses.
The gradient of $L_{\mathrm{VSD}}$ could also be approximated as:
$\nabla_{\theta}L_{\mathrm{VSD}}=\mathbb{E}_{t,\boldsymbol{\epsilon},c}[\omega(t)\|\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t},t,y)-\boldsymbol{\epsilon}_{\text{lora}}(\boldsymbol{x}_{t},t,y,\boldsymbol{c})\|^{2}],$ (2)
where $\epsilon_{lora}$ estimates the denoising direction produced by the customized LoRA model. VSD loss effectively improves the fidelity of both texture and geometry.
Based on VSD, DreamCraft3D [41] proposed bootstrapped score distillation (BSD), which personalizes the generative 3D prior by leveraging multi-view renderings of the 3D scene to finetune the pretrained diffusion model:
$\nabla_{\theta}L_{\mathrm{BSD}}=\mathbb{E}_{t,\boldsymbol{\epsilon},c}[\omega(t)\|\boldsymbol{\epsilon}_{\text{ft}}(\boldsymbol{x}_{t},t,y)-\boldsymbol{\epsilon}_{\text{lora}}(\boldsymbol{x}_{t},t,y,\boldsymbol{c})\|^{2}],$ (3)
where $\epsilon_{\text{ft}}$ denotes the finetuned model.
However, the customization for each text-conditioned 3D scene or image is time-consuming. As for optimizing 3D heads, we directly apply a pretrained ID-aware diffusion model as a strong 3D guidance for in-the-wild human heads.
@@ 4 METHOD
In this paper, we propose Portrait3D, a novel optimization-based pipeline for high-fidelity 360-degree head generation from a single portrait image. We fully leverage state-of-the-art identity-aware priors to our generation process through ID-aware Initialization, ID-aware Guidance, and ID-aware Geometry Supervision4.1). The overall pipeline of Portrait3D is shown in Figure 3, which consists of two stages: geometry sculpting stage and texture generation stage.
In the geometry sculpting stage, we utilize DMTet [40] as the 3D representation and propose ID-aware Score Distillation (ISD) loss to inject identity guidance for the optimization (§4.2). Then in the texture generation stage, we propose ID-Consistent Texture Inpainting and Refinement for realistic appearance generation (§4.3).
@@ 4.1 ID-aware Initialization and Guidance
Given a facial image $\boldsymbol{x}_{\mathrm{ref}}$ as input, our goal is to generate a high-fidelity 3D head model parameterized with $\theta$, that preserves the identity and appearance of the person in the image. We adopt an optimization-based pipeline, and formulate the optimization process for generating 3D head as follows:
$\begin{split}&\min\begin{cases}\mathbb{E}_{t,c}[\operatorname{D}_{\mathrm{KL}}(q_{t}^{\theta}(\boldsymbol{x}_{t}|\mathbf{c})\|p_{t}(\boldsymbol{x}_{t}|y))],&\mathbf{c}=\mathbf{c}_{\text{rand}}\\ D(\boldsymbol{x}_{\mathrm{ref}},\;g(\theta,\boldsymbol{c})),&\mathbf{c}=\mathbf{c}_{\text{ref}}\end{cases}\\ &\text{s.t. }\theta=\{\theta|\boldsymbol{x}_{\mathrm{ref}},\;\theta_{0}\},\quad\theta_{0}=\text{Initial Mesh}\end{split}$ (4)
where $\mathbf{c}_{\text{rand}}$ is a random camera pose, and $\mathbf{c}_{\text{ref}}$ is the reference camera pose (in $\boldsymbol{x}_{\mathrm{ref}}$). $D(\cdot)$ denotes the distance function, and $\theta_{0}$ is the initial 3D mesh provided by users, which is usually an ellipsoid or a parametric model.
For 3D full head generation, the quality of the generated 3D head largely relies on the quality of the initial 3D model and the multi-view guidance. Some existing 3D generation methods use text prompts $y$ that accurately describe the facial image or view-dependent diffusion models [13, 16, 23, 25] to provide additional guidance of the image. However, for one-shot 3D head generation tasks, text prompts can only express limited information:
The direct description of the input facial image, often cannot fully capture the variations of different facial features. Besides, pretrained 2D view-dependent diffusion models (e.g., Zero-123 [23]) are trained on general objects, which often lack the ability to generalize to human heads.
Therefore, our key insight is to fully leverage identity knowledge and inject state-of-the-art identity priors into the mesh initialization and diffusion guidance.
ID-aware Initialization. The initialization is crucial for 3D generation. According to the optimization goal in Eq.(4), a good initial 3D mesh $\theta_{0}$ should possess the approximate geometric shape of the face in the given image $\boldsymbol{x}_{\mathrm{ref}}$. Some approach uses FLAME [8] to obtain the initial 3D head shape. However, this approach cannot generate personalized facial features and does not describe the geometric shape of hair. To address this issue, we employ a 3D-aware GAN model [1] trained on a comprehensive $360^{\circ}$ head dataset to obtain the initial 3D head mesh $\theta_{0}$. Specifically, given the input facial image $\boldsymbol{x}_{\mathrm{ref}}$, 1) we first perform inversion by optimization based on pixel-wise $\mathcal{L}_{2}$ loss and image-level LPIPS loss [50], to find the corresponding latent code $z$; 2) and then perform Pivotal Tuning Inversion (PTI) [36], which uses the initial inverted latent code $z$ as a pivot and finetunes the generator, to obtain a 3D representation that better matches the input image. Through 3D GAN inversion, we can obtain an initial 3D head mesh that corresponds to the input facial image, denoted as:
@@
ID-aware Guidance. In order to obtain finer information about the facial image, we extract high-dimensional identity features $I_{r}$ from the reference image as additional conditions beyond text prompt, to guide the denoising process of diffusion models. We use a facial recognition model [6] trained on large-scale facial datasets as the identity encoder to extract robust identity features. Additionally, we utilize a light projection network to map the extracted ID feature to an enhanced ID feature $f_{ID}$ aligned with text embedding, and insert the ID feature into diffusion with decoupled cross-attention [47]. In this way, we construct an ID-guided diffusion model $p^{\text{id}}$, which will be utilized in our ID-aware SDS loss to provide identity guidance for a specific person.
The aforementioned identity-enhanced prompt only extracts general identity information, but cannot provide the information of facial layout. For example, a person has different mouth shapes when laughing and crying, but the identity features are the same. Therefore, solely using the identity feature-based guidance may lead to inconsistency in the facial layout when viewed from different camera viewpoints. To address this issue, we utilize a pretrained facial shape predictor [11] to obtain 3D facial landmarks of the reference image $\boldsymbol{x}_{\mathrm{ref}}$, and apply pose alignment to align the 3D facial landmarks to the initial 3D head mesh. Then, given a camera pose $\boldsymbol{c}$, we project the extracted 3D facial landmarks onto the 2D plane, and obtain a facial landmark image $L(\mathbf{c})$, which provides the facial component layout information from that viewpoint. We then apply a landmark-guided ControlNet [49] to introduce the facial landmarks information into our ID-aware SDS loss. Since the facial landmark images under different viewpoints are projected from the same 3D facial landmarks, they are geometrically consistent. Incorporating these facial landmark images in our ID-aware SDS will provide multi-view geometry consistency constraints and guidance.
By incorporating three types of information, i.e., text prompt $y$ (obtained from $\boldsymbol{x}_{\mathrm{ref}}$ by BLIP-2 [20]), identity feature $f_{ID}$, and the projected facial landmark image $L(\mathbf{c})$, we propose ID-aware Score Distillation (ISD) loss, which can be formulated as:
@@
ID-aware Geometry Supervision. In the geometry sculpting stage, the reference image $\boldsymbol{x}_{\mathrm{ref}}$ in RGB space cannot be directly used as the supervision for geometry. A common method is to use normal maps and depth maps that contain geometric information for supervision, which can be predicted based on the reference image in RGB space. Therefore, it is crucial to obtain high-quality geometric representations $\boldsymbol{X}_{\text{ref}}^{\text{geo}}$ from the reference image in RGB space. Previous methods [7] directly predicted depth maps and normal maps from images but failed to capture some geometric details in the reference images, which could result in flat geometries, as shown in Fig 4. Different from previous methods, we first use an edge extractor to extract high-frequency facial geometric details and obtain a Canny image as the condition. Then we utilize pre-trained Text-to-Normal and Text-to-Depth diffusion models [16] with ControlNet to generate normal maps and depth maps. Additionally, to better preserve identity information, we also add ID features into the diffusion process. Through these approaches, high-quality geometric supervision in the reference view can be formulated as:
@@
Figure 4: Comparison between our normal map and depth map and those generated by DPT[35, 34].

### v2

# ID-Sculpt: ID-aware 3D Head Generation from Single In-the-wild Portrait Image

@@ Abstract
While recent works have achieved great success on image-to-3D object generation, high quality and fidelity 3D head generation from a single image remains a great challenge. Previous text-based methods for generating 3D heads were limited by text descriptions and image-based methods struggled to produce high-quality head geometry. To handle this challenging problem, we propose a novel framework, ID-Sculpt, to generate high-quality 3D heads while preserving their identities. Our work incorporates the identity information of the portrait image into three parts: 1) geometry initialization, 2) geometry sculpting, and 3) texture generation stages. Given a reference portrait image, we first align the identity features with text features to realize ID-aware guidance enhancement, which contains the control signals representing the face information. We then use the canny map, ID features of the portrait image, and a pre-trained text-to-normal/depth diffusion model to generate ID-aware geometry supervision, and 3D-GAN inversion is employed to generate ID-aware geometry initialization. Furthermore, with the ability to inject identity information into 3D head generation, we use ID-aware guidance to calculate ID-aware Score Distillation (ISD) for geometry sculpting. For texture generation, we adopt the ID Consistent Texture Inpainting and Refinement which progressively expands the view for texture inpainting to obtain an initialization UV texture map. We then use the ID-aware guidance to provide image-level supervision for noisy multi-view images to obtain a refined texture map. Extensive experiments demonstrate that we can generate high-quality 3D heads with accurate geometry and texture from a single in-the-wild portrait image. The project page is at https://jinkun-hao.github.io/ID-Sculpt.
@@ 1 Introduction
Figure 1: ID-Sculpt can generate a high-quality 3D head with detailed face geometry and photo-realistic texture given an in-the-wild portrait image.
@@
3D head generation aims to create high-quality 3D digital assets of human heads that align with user preferences, utilizing various input data, which has significant applications in film character creation, gaming, online meetings, education, etc. Conventional methods rely on multi-view geometric consistency (Kirschstein et al. 2023; Zheng et al. 2023a; Munkberg et al. 2022; Yariv et al. 2021) or statistical 3D face prior model (Bai et al. 2023; Zheng et al. 2023b; Grassal et al. 2022) to achieve high-quality 3D head generation. The former necessitates multi-view images or several minutes of monocular video, which are not applicable to a single in-the-wild portrait image. Moreover, 3D face shape prior models typically focus on the face region, failing to accurately reconstruct the entire head shape and being unable to handle elements such as long hair and clothing. An alternative approach involves employing Generative Adversarial Networks (GANs) to accomplish 3D head generation, which is constrained by limited training data and typically produces frontal-facing faces. Moreover, they struggle to preserve identity consistency when presented with images of large head poses. Recently, text-to-image (Ramesh et al. 2021, 2022; Rombach et al. 2022; Saharia et al. 2022; Han et al. 2023; Pei et al. 2024) generation models trained on large-scale image datasets have achieved significant success. Some approaches employ SDS loss (Poole et al. 2022) to ensure similarity between the generated 3D model and images, facilitating image-to-3D (Melas-Kyriazi et al. 2023; Tang et al. 2023; Liu et al. 2023c, b; Long et al. 2023) or text-to-3D (Poole et al. 2022; Lin et al. 2023; Sanghi et al. 2022; Wang et al. 2022, 2024). However, when employing these methods for customized 3D head generation, the absence of identity information makes it difficult to reproduce the geometric details and texture realism in the portrait image.
@@
Through analyzing previous methods, we summarize the key challenges from the three stages in the optimization-based 3D head generation approach: 1) Geometry Initialization: Previous image-to-3D methods used an ellipsoid as the initial geometry (Chen et al. 2023) and could not accurately reconstruct the geometry of the head. On the other hand, text-to-head methods used a unified FLAME head (Han et al. 2024; Liu et al. 2023a) as the geometric prior, which lacks personalized information and cannot generate individualized features that deviate from the FLAME head, such as long hair. 2) Geometry Sculpting: Although image-to-3D methods attempted to use view-dependent diffusion  (Liu et al. 2023b)to guide geometry generation, the resulting geometry lacked the generalization ability to the reasonable head geometry. Besides, text-to-3D methods use facial image description text to generate geometry, which can result in a deviation from the identity of the reference image. 3) Texture Generation: Previous methods used VSD (Wang et al. 2024) to mitigate color over-saturation issues based on SDS. However, the generated textures still tend to have excessive shading and unrealistic texture (Huang et al. 2023).
To address the challenges mentioned above, we incorporate identity information into the three stages of 3D head generation compared with previous methods in Figure 2. Firstly, we extract Identity embedding using an ID feature extractor and align them with text embedding using a pre-trained ID adapter with decoupled cross-attention. This, along with landmark-guided ControlNet, enables ID diffusion and provides ID-aware guidance, allowing us to incorporate the identity and facial layout information from the reference image into the geometry and texture generation stages. Secondly, we propose ID-aware geometry supervision. By extracting canny maps containing detailed information from the portrait image, we combine canny-guided ControlNet and ID features with off-the-shelf text-to-normal and text-to-depth methods (Huang et al. 2023) to generate high-quality geometric constraint for the geometry generation stage. Additionally, we obtain ID-aware initialization using 3D GAN inversion. In the geometry sculpting stage, we optimize the 3D head using ID-aware guidance with ID-aware Score Distillation (ISD) loss. In the texture generation stage, we first use Progressive Texture Inpainting to generate a UV-textured map for the uncolored 3D head mesh. Then we leverage ID diffusion to denoise the rendered image from random viewpoints and further refine the texture by calculating the loss between the rendered and denoised images.
In this paper, we propose a new method for generating high-quality 3D heads from single in-the-wild portrait images. The main contributions of this paper are as follows:
@@
  • An ID Consistent Texture Inpainting and Refinement method, which uses Progressive Texture Inpainting and refinement method to achieve realistic texture generation.
2 Related Works
2.1 Text to 3D Generation
Recent advances in image generative models make it possible to synthesize diverse high-fidelity 3D objects from text prompts. Efforts such as DreamFusion (Poole et al. 2022) explore text-to-3D generation by leveraging a Score Distillation Sampling (SDS) loss shows impressive results. Subsequently, Magic3D (Lin et al. 2023) increases the rendering resolution of generated 3D objects. Fantasia3D (Chen et al. 2023) models the appearance via the BRDF modeling. ProlificDreamer (Wang et al. 2024) proposes Variational Score Distillation (VSD) to mitigate oversaturation problems. Although many works have solved the problem of texture oversaturation to some extent, it is still difficult to generate realistic geometry and texture.
2.2 Image to 3D Generation
Image to 3D generation has also developed rapidly. Zero-1-to-3 (Liu et al. 2023b) first proposes view-dependent diffusion, which focuses on learning camera viewpoints within ]diffusion models, enabling zero-shot novel view synthesis. GeNVS (Chan et al. 2023) leverages pixel-aligned features from input views to add 3D awareness to 2D diffusion models. On the other hand, some optimize-based image-to-3D methods, such as Magic123 (Qian et al. 2023) and Make-it-3D (Tang et al. 2023), incorporate 3D priors from Zero-1-to-3 and employ coarse-to-fine two-stage optimization. Subsequent works (Sun et al. 2023; Yu et al. 2023) further refining and enhancing the quality of generated 3D models. However, these methods are limited by the ability of view-dependent diffusion, which cannot accurately generate the geometry from the image.
@@ 2.2 Image to 3D Generation
Recently, various approaches (Grassal et al. 2022; Gafni et al. 2021; Zheng et al. 2023b; Xu et al. 2023) have explored reconstructing 3D head avatars from monocular face videos using different 3D representations. However, these methods necessitate several minutes of video input and struggle to reconstruct less prominent parts in videos. In addition, methods generating heads from 3D GANs like EG3D (Chan et al. 2022) proposes a tri-plane representation that can efficiently render high-quality facial images. Panohead (An et al. 2023) advances EG3D by enabling full-head 3D face generation. Rodin (Wang et al. 2023) uses diffusion models to generate 3D heads. Nonetheless, they are constrained by the inherent limitations of implicit representations and cannot produce high-quality surface geometry.
Text-based human head generation methods like Headsculpt (Han et al. 2024) and HeadArtist (Liu et al. 2023a) integrate facial landmarks to solve layout issues like the Janus face problem. Humannorm (Huang et al. 2023) refines diffusion models using 3D human body data for better normal and depth guidance. Despite these, texture saturation in these methods can impede the attainment of photo-realistic results. In contrast, our method incorporates identity information in all generation stages, which can generate 3D head with great texture realism and identity consistency.
@@
Figure 3: Overview of our proposed ID-Sculpt framework. Given a reference image, ID-Sculpt leverages 3D GAN inversion for improved identity-related geometry initialization in stage 1. In stage 2, ID-aware guidance and ID-aware geometry supervision are used to intergrate identity information into the geometry sculpting process. In stage 3, ID Consistent Texture Inpainting and Refinement is applied to generate a high-quality head texture, where we first use the inpainting method to generate a rough texture and then use image-level ID-aware supervision for texture refinement. With these methods, we can generate high-quality 3D head models with consistent identities from a single in-the-wild face image.
@@ 3 Method
In this paper, we propose ID-Sculpt, a novel optimization-based method for high-fidelity 3D head generation from a single portrait image. The overall pipeline of ID-Sculpt is shown in Figure 3. We fully leverage identity-aware priors to our generation process through ID-aware Initialization and Guidance3.1). The generation process consists of two stages: Geometry Sculpting3.2) and ID-Consistent Texture Inpainting and Refinement3.3).
@@ 3.1 ID-aware Initialization and Guidance
Given a facial image $\boldsymbol{x}_{\mathrm{ref}}$ as input, our goal is to generate a high-fidelity 3D head model parameterized with $\theta$, that preserves the identity and appearance of the portrait image. We adopt an optimization-based pipeline based on the pretrained text-to-image diffusion model $p_{t}(\boldsymbol{x}_{t}|y)$, and formulate the optimization process for generating a 3D head as follows:
$\begin{split}&\min\begin{cases}\mathbb{E}_{t,c}[\operatorname{D}_{\mathrm{KL}}(q_{t}^{\theta}(\boldsymbol{x}_{t}|\mathbf{c})\|p_{t}(\boldsymbol{x}_{t}|y))],&\mathbf{c}=\mathbf{c}_{\text{rand}}\\ D(\boldsymbol{x}_{\mathrm{ref}},\;g(\theta,\boldsymbol{c})),&\mathbf{c}=\mathbf{c}_{\text{ref}}\end{cases}\\ &\text{s.t. }\theta=\{\theta|\boldsymbol{x}_{\mathrm{ref}},\;\theta_{0}\},\quad\theta_{0}=\text{Initial Mesh},\end{split}$ (1)
where $p_{t}(\boldsymbol{x}_{t}|y)$ is the pretrained text-to-image diffusion model and $q_{t}^{\theta}(\boldsymbol{x}_{t}|\mathbf{c})$ is the distribution at time $t$ of the forward diffusion process starting from the rendered image $g(\theta,\boldsymbol{c})$ with the camera $\mathbf{c}$. $\mathbf{c}_{\text{ref}}$ is the reference camera pose. $D(\cdot)$ denotes the distance function, and $\theta_{0}$ is the initial 3D mesh provided by users.
For 3D full head generation task, the direct description of the input facial image, often cannot fully capture the variations of different facial features. Besides, pretrained 2D view-dependent diffusion models (e.g., Zero-123 (Liu et al. 2023b)) are trained on general objects, which cannot generalize to human heads. Therefore, our key insight is to leverage identity knowledge and inject state-of-the-art identity priors into the mesh initialization and diffusion guidance.
ID-aware Initialization. The initialization is crucial for 3D generation. According to the optimization goal in Eq.(1), a good initial 3D mesh $\theta_{0}$ should possess the approximate geometric shape of the face in the given image $\boldsymbol{x}_{\mathrm{ref}}$. Some approach uses FLAME (Feng et al. 2021) to obtain the initial 3D head shape. However, this approach cannot generate personalized facial features and does not describe the geometric shape of hair. To address this issue, we employ a 3D-aware GAN model (An et al. 2023) trained on a comprehensive $360^{\circ}$ head dataset to obtain the initial 3D head mesh $\theta_{0}$. Specifically, given the input facial image $\boldsymbol{x}_{\mathrm{ref}}$, we first perform inversion to find latent code $z$ and then perform Pivotal Tuning Inversion (PTI) (Roich et al. 2022) to obtain a 3D representation that better matches the input image. Through 3D GAN inversion, we can obtain an initial 3D head mesh that corresponds to the input facial image, denoted as:
@@
ID-aware Guidance. In order to obtain finer information about the facial image, we extract high-dimensional identity features $I_{r}$ from the reference image as additional conditions beyond text prompt, to guide the denoising process of diffusion models. We use a facial recognition model (Deng et al. 2019) trained on large-scale facial datasets as the identity encoder to extract robust identity features. Additionally, we utilize a light projection network to map the extracted ID feature to an enhanced ID feature $f_{ID}$ aligned with text embedding, and insert the ID feature into diffusion with decoupled cross-attention (Ye et al. 2023). In this way, we construct an ID-guided diffusion model $p^{\text{id}}$, which will be utilized in our ID-aware Score Distillation loss to provide identity guidance for a specific person.
Identity-enhanced prompt only extracts general identity information, but cannot provide the information of 3D facial layout. For example, a person has different mouth shapes when laughing and crying, but the identity features are the same. To address this issue, we utilize a pretrained facial shape predictor (Guo et al. 2020) to obtain 3D facial landmarks of the reference image $\boldsymbol{x}_{\mathrm{ref}}$, and align the 3D facial landmarks to the initial 3D head mesh.
Given a camera pose $\boldsymbol{c}$, we project the extracted 3D facial landmarks onto the 2D plane, and obtain a facial landmark image $L(\mathbf{c})$, which provides the facial component layout information from that viewpoint. We then apply a landmark-guided ControlNet (Zhang, Rao, and Agrawala 2023) to introduce the facial layout information into diffusion guidance.
By incorporating three types of information, i.e., text prompt $y$ (obtained from $\boldsymbol{x}_{\mathrm{ref}}$ by BLIP-2 (Li et al. 2023)), identity feature $f_{ID}$, and the projected facial landmark image $L(\mathbf{c})$, we propose ID-aware Score Distillation (ISD) loss, which can be formulated as:
@@
ID-aware Geometry Supervision. In the geometry sculpting stage, the reference image $\boldsymbol{x}_{\mathrm{ref}}$ in RGB space cannot be directly used as the supervision for geometry. A common method is to use normal maps and depth maps that contain geometric information for supervision, which can be predicted based on the reference image in RGB space. Therefore, it is crucial to obtain high-quality geometric representations $\boldsymbol{X}_{\text{ref}}^{\text{geo}}$ from the reference image in RGB space. Previous methods (Eftekhar et al. 2021) directly predicted depth maps and normal maps from images but failed to capture some geometric details in the reference images, which could result in flat geometries, as shown in Fig 4. Different from previous methods, we first use an edge extractor to extract high-frequency facial geometric details and obtain a Canny image as the condition. Then we utilize pre-trained Text-to-Normal and Text-to-Depth diffusion models (Huang et al. 2023) with ControlNet to generate normal maps and depth maps. Additionally, to better preserve identity information, we also add ID features into the diffusion process. Through these approaches, high-quality geometric supervision in the reference view can be formulated as:
@@
Figure 4: Comparison between our normal map and depth map and those generated by DPT(Ranftl, Bochkovskiy, and Koltun 2021; Ranftl et al. 2020).